# General relativistic self-gravitating equilibrium disks around rotating neutron stars

### v1

@@ Abstract
In modeling a relativistic disk around a compact object, the self-gravity of the disk is often neglected while it needs to be incorporated for more accurate descriptions in several circumstances. Extending the Komatsu-Eriguchi-Hachisu self-consistent field method, we present numerical models of a rapidly rotating neutron star with a self-gravitating disk in stationary equilibrium. In particular, our approach allows us to obtain numerical solutions involving a massive disk with the rest mass $\mathcal{O}(10^{-1})-\mathcal{O}(10^{0})\,M_{\odot}$ closely attached to a rotating neutron star. We also assess the impact of self-gravity on the internal structure of the disk and the neutron star. These axisymmetric, stationary solutions can be employed for simulations involving the neutron star-disk system in the context of high-energy transients and gravitational wave emissions.
@@ 1 Introduction
Numerical schemes for studying the structure of rotating neutron stars have been developed in a variety of flavors (see e.g. Paschalidis & Stergioulas (2017) for a review on this topic). Of particular interest is the work by Hachisu (1986), who proposed a novel iterative self-consistent field method to construct equilibrium models of rotating stars in the Newtonian framework. Based on a similar strategy, Komatsu, Eriguchi & Hachisu (1989a) constructed equilibrium models of uniformly rotating neutron stars in full general relativity, and subsequently extended to differential rotation (Komatsu et al., 1989b). This so-called the KEH method, which is named after the three authors of the original work, has been widely employed for studying rotating neutron stars (e.g. Cook et al., 1992, 1994a, 1994b; Eriguchi et al., 1994; Stergioulas & Friedman, 1995; Nozawa et al., 1998; Bauswein & Stergioulas, 2017; Iosif & Stergioulas, 2022a, b) along with some modifications made in later studies.
@@ 2.2.2 Equipotential surfaces and matter distribution
In this work, we assume that the distribution of the specific angular momentum on the equatorial plane follows the power law
@@ 3.1 Computational procedure
Before entering the cycle of iterations (i)-(v) for the first time, we start with a value of $W_{\text{in}}$ very close to the minimum value of $W_{\text{eq}}(r)$ to converge to a solution with a disk of a very small mass first. Then the value of $W_{\text{in}}-W_{\text{center}}$ is increased by a small increment to perform another set of iterations until the solution converges to a new state with a slightly larger disk mass. By repeating this procedure, a sequence of numerical models with increasing disk mass can be obtained.
@@ 4.2.1 Constant angular momentum disk
(a) $\kappa=5.31$, $M_{\mathrm{Disk}}=1.00\times 10^{-3}$
(b) $\kappa=5.52$, $M_{\mathrm{Disk}}=1.67\times 10^{-2}$
(c) $\kappa=6.20$, $M_{\mathrm{Disk}}=5.97\times 10^{-2}$
(a) $\kappa=6.62$, $M_{\mathrm{Disk}}=5.66\times 10^{-3}$
(b) $\kappa=7.40$, $M_{\mathrm{Disk}}=0.760$
(c) $\kappa=7.88$, $M_{\mathrm{Disk}}=1.02$
Figure 5: Same as Figure 4, for rapidly rotating case $r_{p}/r_{e}=0.60$.
(a) $\kappa=6.95$, $M_{\mathrm{Disk}}=7.40\times 10^{-3}$
(b) $\kappa=7.48$, $M_{\mathrm{Disk}}=0.419$
(c) $\kappa=8.22$, $M_{\mathrm{Disk}}=0.242$
@@
Models D1-D4 have a rapidly rotating neutron star with a higher maximum rest energy density $\rho_{0}^{\mathrm{max}}=2.0\times 10^{-3}$; models D1, D2, and D4 are shown in Figure 6. The neutron star has a higher mass $M_{\mathrm{NS}}\approx 2.0$. more compact radius $r_{e}\approx 14\text{km}$, and shorter rotation period $\leq 1\text{ms}$. This sequence yields models with the disk mass being a bit smaller than the models B1-B4.
@@ 4.2.2 Non-constant angular momentum disk
(a) $\kappa=5.39$, $M_{\mathrm{Disk}}=1.68\times 10^{-5}$
(b) $\kappa=5.42$, $M_{\mathrm{Disk}}=2.44\times 10^{-4}$
(c) $\kappa=5.51$, $M_{\mathrm{Disk}}=7.09\times 10^{-5}$
(a) $\kappa=5.42$, $M_{\mathrm{Disk}}=1.05\times 10^{-7}$
(b) $\kappa=5.44$, $M_{\mathrm{Disk}}=1.46\times 10^{-5}$
(c) $\kappa=5.47$, $M_{\mathrm{Disk}}=6.68\times 10^{-5}$
@@ 4.3 Comparisons with non-self-gravitating disks
Figure 9: Comparison between models with and without the disk self-gravity for $\kappa=7.88$, $M_{\mathrm{Disk}}=0.101$. (Top) Rest mass density of the disk in self-gravitating and non-self-gravitating cases. To better visualize the difference of the matter distribution between two cases, we draw contour lines for $\rho_{0}=10^{-9},10^{-8},10^{-7}$ with solid, dotted, and dashed white lines, respectively. (Bottom) Effective potential $W_{\text{eq}}(r)$. The inner and outer edges of the disk are shown with small circles, and the interior region of the disk is displayed with colored shades.
Figure 10: Same as Figure 9, with $\kappa=7.78$ and $M_{\mathrm{Disk}}=0.969$. In the upper panel, an additional contour level $\rho_{0}=10^{-6}$ is shown with a white dash-dotted line.
@@
Figure 12: Rest mass of the disk versus the depth parameter $w$. Non self-gravitating disk solutions are shown with dashed lines while the self-gravitating solutions are shown with solid lines. Model parameters are $\rho_{0}^{\mathrm{max}}=1.0\times 10^{-3}$, $r_{p}/r_{e}=0.60$, and $a=0.0$. A kink of the $\kappa=7.78$ curve near $M_{\mathrm{Disk}}=0.8$ is a parametrization artifact; see discussions in Sec. 4.3.
@@ 5 Summary and conclusion
The uniformly rotating assumption on the neutron star greatly simplifies the problem, but at the same time largely restricting us from resolving radius-dependent effects by the gravitational pull of the disk to the interior of the neutron star. Further research adopting a more realistic, differential rotation law for the neutron star (e.g. Galeazzi et al., 2012; Uryū et al., 2017; Bauswein & Stergioulas, 2017; Bozzola et al., 2018; Iosif & Stergioulas, 2022a) will allow us to examine more diverse configurations of the neutron star-disk system.
@@
Since our models do not take energy and pressure contributions from magnetic fields into account, they are only applicable to an unmagnetized or at most weakly magnetized system. In realistic scenarios, magnetic fields would trigger instabilities so that such systems evolve into accreting states. The dynamical stability of our results is still under question as well. These aspects can be further investigated by feeding our model to an evolution code as initial data and running general relativistic hydrodynamics simulations, which we aim as future work.
@@ Acknowledgements
YK thanks Chan Park for his assistance and helpful comments. JK acknowledges that this research was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1I1A2050775). HIK acknowledges that this research was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education through the Center for Quantum Spacetime (CQUeST) of Sogang University (NRF-2020R1A6A1A03047877). HML was supported by the NRF grant No. 2021M3F7A1082056. We acknowledge the hospitality at APCTP where part of this work was done. Computations were performed on the Gmunu cluster at the Korea Astronomy and Space Science Institute. Figures in this article were produced using Matplotlib (Caswell et al., 2023), Numpy (Harris et al., 2020), and Scipy (Virtanen et al., 2020) packages.

### v2

@@ Abstract
In modeling a relativistic disk around a compact object, the self-gravity of the disk is often neglected while it needs to be incorporated for more accurate descriptions in several circumstances. Extending the Komatsu-Eriguchi-Hachisu self-consistent field method, we present numerical models of a rapidly rotating neutron star with a self-gravitating disk in stationary equilibrium. In particular, our approach allows us to obtain numerical solutions involving a massive disk with the rest mass $\mathcal{O}(10^{-1})-\mathcal{O}(10^{0})\,M_{\odot}$ closely attached to a rotating neutron star, given that the disk is mainly supported by the relativistic electron degeneracy pressure. We also assess the impact of self-gravity on the internal structure of the disk and the neutron star. These axisymmetric, stationary solutions can be employed for simulations involving the neutron star-disk system in the context of high-energy transients and gravitational wave emissions.
@@ 1 Introduction
Numerical schemes for studying the structure of rotating neutron stars have been developed in a variety of flavors (see e.g. Paschalidis & Stergioulas (2017) for a review on this topic). Of particular interest is the work by Hachisu (1986), who proposed a novel iterative self-consistent field method to construct equilibrium models of rotating stars in the Newtonian framework. Based on a similar strategy, Komatsu, Eriguchi & Hachisu (1989a) constructed equilibrium models of uniformly rotating neutron stars in full general relativity, and subsequently extended to differential rotation (Komatsu et al., 1989b). This so-called the KEH method, which is named after the three authors of the original work, has been widely employed for studying rotating neutron stars (e.g. Cook et al., 1992, 1994a, 1994b; Eriguchi et al., 1994; Stergioulas & Friedman, 1995; Nozawa et al., 1998; Bauswein & Stergioulas, 2017; Iosif & Stergioulas, 2021, 2022) along with some modifications made in later studies.
@@ 2.2.2 Equipotential surfaces and matter distribution
In this work, we assume that the distribution of the specific angular momentum on the equatorial plane follows the power law (e.g. Daigne & Font, 2004)
@@ 3.1 Computational procedure
Before entering the cycle of iterations (i)-(v) for the first time, we start with a value of $W_{\text{in}}$ very close to the minimum value of $W_{\text{eq}}(r)$ to converge to a solution with a disk of a very small mass first. Then the value of $W_{\text{in}}-W_{\text{center}}$ is increased by a small increment to perform another set of iterations until the solution converges to a new state with a slightly larger disk mass. In other words, we start with a light disk and use it as an initial guess (seed solution) for a heavier disk progressively. By repeating this procedure, a sequence of numerical models with increasing disk mass can be obtained. Our strategy outlined above requires multiple rounds of iterations to reach a model with a massive disk, starting from and iterating through the ones with a lighter disk. While directly approaching to a single point chosen in the parameter space can be seemingly computationally more efficient, we have empirically observed that if the resulting disk is too massive, iterations become quickly unstable and fail within few cycles.
@@ 4.2 Equilibrium Models
(a) $\kappa=5.31$, $M_{\mathrm{Disk}}=1.00\times 10^{-3}M_{\odot}$
(b) $\kappa=5.52$, $M_{\mathrm{Disk}}=1.67\times 10^{-2}M_{\odot}$
(c) $\kappa=6.20$, $M_{\mathrm{Disk}}=5.97\times 10^{-2}M_{\odot}$
(a) $\kappa=6.62$, $M_{\mathrm{Disk}}=5.66\times 10^{-3}M_{\odot}$
(b) $\kappa=7.40$, $M_{\mathrm{Disk}}=0.760M_{\odot}$
(c) $\kappa=7.88$, $M_{\mathrm{Disk}}=1.02M_{\odot}$
Figure 5: Same as Figure 4, for rapidly rotating case $r_{p}/r_{e}=0.60$.
(a) $\kappa=6.95$, $M_{\mathrm{Disk}}=7.40\times 10^{-3}M_{\odot}$
(b) $\kappa=7.48$, $M_{\mathrm{Disk}}=0.419M_{\odot}$
(c) $\kappa=8.22$, $M_{\mathrm{Disk}}=0.242M_{\odot}$
@@ 4.2.1 Constant angular momentum disk
Models D1-D4 have a rapidly rotating neutron star with a higher maximum rest energy density $\rho_{0}^{\mathrm{max}}=2.0\times 10^{-3}$; models D1, D2, and D4 are shown in Figure 6. The neutron star has a higher mass $M_{\mathrm{NS}}\approx 2.0M_{\odot}$. more compact radius $r_{e}\approx 14\text{km}$, and shorter rotation period $\leq 1\text{ms}$. This sequence yields models with the disk mass being a bit smaller than the models B1-B4.
(a) $\kappa=5.39$, $M_{\mathrm{Disk}}=1.68\times 10^{-5}M_{\odot}$
(b) $\kappa=5.42$, $M_{\mathrm{Disk}}=2.44\times 10^{-4}M_{\odot}$
(c) $\kappa=5.51$, $M_{\mathrm{Disk}}=7.09\times 10^{-5}M_{\odot}$
(a) $\kappa=5.42$, $M_{\mathrm{Disk}}=1.05\times 10^{-7}M_{\odot}$
(b) $\kappa=5.44$, $M_{\mathrm{Disk}}=1.46\times 10^{-5}M_{\odot}$
(c) $\kappa=5.47$, $M_{\mathrm{Disk}}=6.68\times 10^{-5}M_{\odot}$
@@ 4.3 Comparisons with non-self-gravitating disks
Figure 9: Comparison between models with and without the disk self-gravity for $\kappa=7.88$, $M_{\mathrm{Disk}}=0.101M_{\odot}$. (Top) Rest mass density of the disk in self-gravitating and non-self-gravitating cases. To better visualize the difference of the matter distribution between two cases, we draw contour lines for $\rho_{0}=10^{-9},10^{-8},10^{-7}$ with solid, dotted, and dashed white lines, respectively. (Bottom) Effective potential $W_{\text{eq}}(r)$. The inner and outer edges of the disk are shown with small circles, and the interior region of the disk is displayed with colored shades.
Figure 10: Same as Figure 9, with $\kappa=7.78$ and $M_{\mathrm{Disk}}=0.969M_{\odot}$. In the upper panel, an additional contour level $\rho_{0}=10^{-6}$ is shown with a white dash-dotted line.
@@
Figure 12: Rest mass of the disk (in units of the solar mass) versus the depth parameter $w$. Non self-gravitating disk solutions are shown with dashed lines while the self-gravitating solutions are shown with solid lines. Model parameters are $\rho_{0}^{\mathrm{max}}=1.0\times 10^{-3}$, $r_{p}/r_{e}=0.60$, and $a=0.0$. A kink of the $\kappa=7.78$ curve near $M_{\mathrm{Disk}}=0.8M_{\odot}$ is a parametrization artifact; see discussions in Sec. 4.3.
@@ 5 Summary and conclusion
The uniformly rotating assumption on the neutron star greatly simplifies the problem, but at the same time largely restricting us from resolving radius-dependent effects by the gravitational pull of the disk to the interior of the neutron star. A more realistic, differential rotation law (e.g. Galeazzi et al., 2012; Uryū et al., 2017; Bauswein & Stergioulas, 2017; Bozzola et al., 2018; Iosif & Stergioulas, 2021) can be used for an improved modeling of the rotating neutron star. Also, we only have considered a simple power law (Eq. (38)) as a representative example of a non-constant angular momentum distribution of the disk, where several alternate prescriptions are available, such as adopting a different form of angular momentum distribution (Penna et al., 2013; Qian et al., 2009) or specifying a nontrivial functional relationship between the angular velocity $\Omega$ and the specific angular momentum $l$ (Witzany & Jefremov, 2018; Wielgus et al., 2015). Further research with an extended set of rotation profiles of both the neutron star and the disk will allow us to examine more diverse configurations of the system.
Several potential improvements for speeding up the computational procedure (described in Sec. 3.1) can be available, such as an adaptive control of the convergence threshold $\Delta$ and the increment in the disk depth parameter $w$ when approaching a solution, or applying the successive over-relaxation method (Varga, 2000; Hageman & Young, 2004) over each iterations.
@@
Since our models do not take energy and pressure contributions from magnetic fields into account, they are only applicable to an unmagnetized or at most weakly magnetized system. In realistic scenarios, magnetic fields can change the equilibrium configuration (e.g. Komissarov, 2006) or would trigger instabilities so that such systems evolve into accreting states. The dynamical stability of our results is still under question as well. These aspects can be further investigated by feeding our model to an evolution code as initial data and running general relativistic hydrodynamics simulations, which we aim as future work.
@@ Acknowledgements
YK thanks to Chan Park for his assistance and comments, and is grateful to Isaac Legred and Michael Pajkos for helpful discussions. JK acknowledges that this research was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education (NRF-2021R1I1A2050775). HIK acknowledges that this research was supported by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education through the Center for Quantum Spacetime (CQUeST) of Sogang University (NRF-2020R1A6A1A03047877). HML was supported by the NRF grant No. 2021M3F7A1082056. We acknowledge the hospitality at APCTP where part of this work was done. Computations were performed on the gmunu cluster at the Korea Astronomy and Space Science Institute. Figures in this article were produced using Matplotlib (Caswell et al., 2023), Numpy (Harris et al., 2020), and Scipy (Virtanen et al., 2020) packages. We also thank the anonymous referee for valuable comments and insightful suggestions.
Data Availability
The computed models which are listed in Table 4 and assigned with Figures (i.e. all entries except A4, B4, D3) are publicly available on Zenodo (Kim et al., 2024). Other data underlying this article will be shared upon request to the corresponding author.